# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

### v2

@@ 4 Probabilistic programming
Using atomic coordinate data, all molecules with stereoisomerism may be represented and distinguished, and dihedral bond angles, torsion angles, and bond lengths may be calculated. Such information adds minimally to the space complexity, but yet can be crucial when predicting physical propreties of drugs.
@@ 5.1 Haskell
What is novel is the integration of probabilistic programming with Haskell into a concrete use case. Probabilistic programming languages themselves in haskell have been created before[54, 55, 42], laying the groundwork for usage with chemical structures. The existence of libraries such as LazyPPL[56] (demonstrated later in the paper), makes the ADT possible to use with Bayesian machine learning applications, such as drug design, which have been attempted before [57, 58, 54], and is an emerging area of research. The “lazy” (non strict, or call-by-need) property of Haskell, as discussed earlier, can provide efficiency, for instance when sampling from expensive functions or performing simulations only when necessary.
The record syntax in Haskell also enables access of only parts of the molecule needed for a task , e.g. in machine learning. Non-necessary data is not accessed, making machine learning models potentially less wasteful. This is difficult to do with non-local encoding of information, such as with SMILES or SELFIES strings, where a whole molecule, as a string, must be parsed.
@@
Furthermore, we have dihedral and bond angles already. We have the meta characters as operations in haskell such as folds for a for loop and filter or takeWhile for a while loop.
@@ 5.2 Record Syntax
Record syntax allows for defining named fields within data types, facilitating direct access to individual parts of a data structure. The use of these names acts as a way to prevent confusion about whether one has a specific atom or coordinate in a data structure representing atoms, for instance. Here is a simple example demonstrating record syntax:
1data Atom = AtomS { atomID :: Integer,
2 symbol :: AtomicSymbol }
3 | AtomC { atomID :: Integer,
4 symbol :: AtomicSymbol,
@@
7data AtomicSymbol = C | O | H
8
9data Coordinate = Coordinate { x :: Double,
@@
Listing 2: Updating the coordinate of an atom
In this example, Atom is a sum type with two constructors: AtomS and AtomC. The AtomS constructor represents an atom with a symbol and has two fields: atomID of type Integer and symbol of type AtomicSymbol. The AtomC constructor represents an atom with a symbol and a coordinate, and it has three fields: atomID of type Integer, symbol of type AtomicSymbol, and coordinate of type Coordinate. Record syntax automatically generates getter functions for each field.
Example 1: Updating the x-coordinate of an atom
@@ 3 AtomC atomID symbol coordinate { x = newX }
Listing 3: Modifying only the x-coordinate of an atom
The setCoordinate function updates the coordinate of an AtomC. It takes a new Coordinate and an Atom as arguments. If the atom is an AtomC, it uses the @ pattern to bind the entire AtomC value to the variable atomWC. This allows us to refer to the original AtomC value when updating its fields. We then use record syntax to create a new AtomC by updating only the coordinate field of atomWC with the newCoordinate, while keeping the atomID and symbol fields unchanged. If the atom is an AtomS, it returns the original atom unchanged since AtomS does not have a coordinate field.
Example 2: Extracting the x-coordinate of an atom
@@
Listing 4: Retrieving the x-coordinate of an atom
The getXCoordinate function retrieves the x-coordinate of an AtomC. It uses pattern matching to extract the coordinate field from the AtomC constructor. If the atom is an AtomC, it uses record syntax to access the x field of the coordinate and returns it wrapped in Just. If the atom is an AtomS, it returns Nothing since AtomS does not have a coordinate field.
@@ 5.3 The Molecule ADT
The following presents the heart of this paper, the molecule data type:
1
2data Molecule = Molecule {
@@
6
7data Atom = Atom {
@@
13
14data BondType = Bond {delocNum :: Integer, atomIDs :: Maybe [Integer]}
@@
19
20data ElementAttributes = ElementAttributes
@@
25
26type EquilibriumBondLength = Angstrom
27newtype Angstrom = Angstrom Double deriving (Read, Show, Eq)
Listing 5: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent Bonds (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over).
@@
ADTs are composed of simple, atomic data types combined using constructs such as products (tuples or records) and sums (variants or unions).
@@ 5.3.1 Example of a Molecule : Benzene
1
2benzene :: Molecule
3benzene = Molecule
4 { atoms = [ Atom i (elementAttributes e) (Coordinate x y z)
@@
7 }
8
9atomsData :: [(Int, Element, Double, Double, Double)]
10atomsData =
11 [ (1, C, 0.0, 1.3948, 0.0), (2, C, 1.2079, 0.6974, 0.0),
12 (3, C, 1.2079, -0.6974, 0.0), (4, C, 0.0, -1.3948, 0.0),
13 (5, C, -1.2079, -0.6974, 0.0), (6, C, -1.2079, 0.6974, 0.0),
14 (7, H, 0.0, 2.4732, 0.0), (8, H, 2.1431, 1.2366, 0.0),
15 (9, H, 2.1431, -1.2366, 0.0), (10, H, 0.0, -2.4732, 0.0),
16 (11, H, -2.1431, -1.2366, 0.0), (12, H, -2.1431, 1.2366, 0.0)
17 ]
18
19bondsData :: [((Int, Int), Bond)]
20bondsData =
@@
22 [ ((1,2), single), ((1,6), single), ((2,3), single),
@@
Listing 6: A representation of the molecule benzene expressed in the Molecule data type, written as a .hs file, and using list comprehension to ensure readability, excluding the Shells currently.
@@
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as Bond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6], explicitly specifying the cyclic structure formed by the carbon atoms.
@@ 5.4 The Orbital ADT
The representation captures quantum mechanical properties of atoms and can be made to adhere to the principles governing the arrangement of electrons in an atom. Each shell is represented as a distinct entity, characterized by its principal quantum number. This aligns with the quantum mechanical description of atomic structure, where electrons occupy discrete energy levels or shells. By organizing the electronic configuration into separate shells, the representation accurately reflects the hierarchical nature of atomic orbitals and the increasing energy and complexity of the electronic structure as the principal quantum number increases.
@@
The use of a list to represent the shells implies that there could be an arbitrarily large finite number of shells (or infinite), which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
@@
Within each shell, there are subshells ($s,p,d$, and $f$), which correspond to the different types of atomic orbitals. Any interface to this data structure must respect the maximum number of electrons that can occupy each subshell, with $s$ subshells containing up to two electrons, $p$ subshells containing up to six electrons, $d$ subshells containing up to ten electrons, and $f$ subshells containing up to fourteen electrons. However, since the current implementation does not utilize dependent types, the correctness of the number of electrons in each subshell cannot be guaranteed at compile time and is left as future work. The responsibility of ensuring the proper electronic configuration lies with the user of the ADT.
1-- Orbital types for each subshell
@@
3data P = Px | Py | Pz deriving (Show, Eq, Read)
4data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read)
5data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
6
7-- The Orbital has a type and a number of electrons.
8data Orbital subshellType = Orbital {
@@
13
14-- Each SubShell consists of a list of Orbitals.
15-- We ensure that all orbitals within a subshell are of the same type
16-- by using the subshellType parameter.
17newtype SubShell subshellType = SubShell {
@@
20
21-- Each Shell has a principal quantum number and consists of a list of SubShells.
22data Shell = Shell {
@@
Listing 7: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells.
Each atomic orbital, contains the subshell type, electron count, and spatial orientation of the orbital. These properties can be used to construct the actual basis functions, but they are not the basis functions themselves. Furthermore, the representation allows for the specification of the spatial orientation of each orbital within a subshell through the use of the Coordinate type. This information is crucial for describing the shape and symmetry of atomic orbitals, particularly for $p$, $d$, and $f$ orbitals that have distinct orientations. However, it is important to note that the actual spatial representation of the orbitals would require the use of the appropriate mathematical functions, such as Slater-type orbitals (STOs), spherical harmonics, or Gaussian functions. Slater-type orbitals are commonly used in quantum chemistry as they provide a more accurate representation of the electron density near the nucleus compared to Gaussian functions. However, these mathematical functions are not explicitly included in this ADT system, which focuses on the symbolic representation of the electronic configuration.
By designing the appropriate interface functions that operate on the Shells type, one can ensure that rules such as the Aufbau principle, Hund’s rule, and the Pauli exclusion principle are properly enforced when manipulating the electronic configuration. These interface functions would act as a layer of abstraction, encapsulating the logic and rules governing the distribution of electrons within the atom. By providing a well-defined set of operations that adhere to these principles, the interface can maintain the integrity and consistency of the electronic configuration, thereby accurately representing the charge of the molecule. The interface functions would handle the creation, modification, and querying of the Shells data structure, ensuring that any changes made to the electronic configuration are valid and compliant with the quantum mechanical principles. This approach allows for a clear separation of concerns, where the Shells type focuses on the data representation, while the interface functions handle the logic and enforcement of the principles governing the electronic configuration.
The Aufbau principle can be enforced by ensuring that electrons are added to orbitals in the correct order of increasing energy. Hund’s rule can be implemented by maximizing the number of unpaired electrons in degenerate orbitals before pairing them. The Pauli exclusion principle can be enforced by limiting the maximum number of electrons in each orbital to two, with opposite spins. The use of type information can greatly assist in enforcing these principles and ensuring the validity of the electronic configuration. For example, the SubShell data type could be parameterized by the maximum number of orbitals allowed for each subshell type. This type-level information would prevent the creation of invalid subshells with more orbitals than permitted by the quantum mechanical rules. Similarly, the electronCount field of the Orbital data type can be constrained to a range of values (e.g., 0 to 2) using type-level naturals or refined types. This ensures that the Pauli exclusion principle is upheld, and no orbital contains more than two electrons.
@@
However, it does not explicitly account for the spin of electrons or the coupling between orbital and spin angular momenta, which play a significant role in determining the detailed electronic structure of atoms, particularly for elements with partially filled subshells.
@@
. Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbitals with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
@@
As mentioned, in Haskell, algebraic data types (ADTs) are constructed using sum and product type constructors, analogous to addition and multiplication in algebra. A sum type is a choice between different types (denoted by the | symbol), akin to an or; for example, a Maybe type represents a value that may be Nothing (like zero in addition) or Just a where a is any type. This is reflective of a disjoint sum in algebra, where a value can be one thing or another, but not both simultaneously. This allows the grammar to distinguish between empty values and zero values. As will be discussed, the orientation field of the Orbital data type is wrapped in a Maybe type, which allows for the natural expression of whether an orbital has a specific orientation or not. For $s$ orbitals, the orientation is Nothing because $s$ orbitals are spherically symmetric and have no spatial orientation. In contrast, $p$ orbitals have specific orientations, and their orientation can be represented as Just Coordinate, where Coordinate is a type that can represent orientation in the $x,y,z$ axes. On the other hand, a product type combines several values into one composite type, similar to how factors multiply to a product; a tuple (a, b) is an instance of a product type containing both a and b.

### v3

@@ 2.1 Our Algebraic Data Type
Using atomic coordinate data, all molecules with stereoisomerism may be represented and distinguished, and dihedral bond angles, torsion angles, and bond lengths may be calculated. Such information adds minimally to the space complexity, but yet can be crucial when predicting physical propreties of drugs. In our ADT, we have dihedral and bond angles. We also have the ‘meta’ characters as operations in Haskell ([1]) such as folds, map and forM_ that together can capture the essence of a for loop and filter or takeWhile for a while loop (perhaps in conjunction with a State monad).
@@ 2.2 Haskell
What is novel is the integration of probabilistic programming with Haskell into a concrete use case. Probabilistic programming languages themselves in haskell have been created before [60, 48, 47, 49], laying the groundwork for usage with chemical structures. The existence of libraries such as LazyPPL[61] (demonstrated later in the article), makes the ADT possible to use with Bayesian machine learning applications, such as drug design, which have been attempted before [62, 52, 60], and is an emerging area of research. The “lazy” (non strict, or call-by-need) property of Haskell, as discussed earlier, can provide efficiency, for instance when sampling from expensive functions or performing simulations only when necessary. Furthermore, it can represent arbitrarily large finite data structures, which sits between an extensional notion of infinite and finiteness. This is useful later on as [1] quotes Generally, a graph with an infinite number of edges cannot be described by a string of finite length. Fortunately, a net can be represented by a finite graph, known as its quotient graph.
@@
Record syntax in Haskell also enables the access of only parts of the molecule needed for a task, e.g. only taking the atomicSymbol of the atoms in a molecule, for a machine learning task. Non-necessary data is not accessed, making machine learning models potentially less wasteful. This is difficult to do with non-local encoding of information, such as with SMILES or SELFIES strings, where a whole molecule, as a string, must be parsed.
Definition 10 (Record Types in Haskell).
A record type is a Haskell data type whose constructors include named fields, enabling direct access to components via automatically generated selector functions.
@@ 2.3 Record Syntax
Record syntax allows for defining named fields within data types, facilitating direct access to individual parts of a data structure. The use of these names acts as a way to prevent confusion about whether one has a specific atom or coordinate in a data structure representing atoms, for instance.
Listing 4 shows a simple example. Atom is a sum (coproduct) type with two constructors: AtomS and AtomC. The AtomS constructor represents an atom with a symbol and has two fields: atomID of type Integer and symbol of type AtomicSymbol. The AtomC constructor represents an atom with a symbol and a coordinate, and it has three fields: atomID of type Integer, symbol of type AtomicSymbol, and coordinate of type Coordinate.
1data Atom = AtomS { symbol :: AtomicSymbol }
2 | AtomC { symbol :: AtomicSymbol,
@@
7data Coordinate = Coordinate { x :: Double,
@@
Listing 4: Haskell record syntax showing a simple algebraic data type
Record syntax automatically generates getter functions for each field. Listing 5 shows the updateXCoordinate function which updates the coordinate of an AtomC. It takes a new Coordinate and an Atom as arguments. If the atom is an AtomC, it uses the @ pattern to bind the entire AtomC value to the variable atomWC. This allows us to refer to the original AtomC value when updating its fields. We then use record syntax to create a new AtomC by updating only the coordinate field of atomWC with the newCoordinate, while keeping the atomID and symbol fields unchanged. If the atom is an AtomS, it returns the original atom unchanged since AtomS does not have a coordinate field.
@@ 3 AtomC atomID symbol coordinate { x = newX }
Listing 5: Modifying only the x-coordinate of an atom
The getXCoordinate function shown in Listing 6 retrieves the x-coordinate of an AtomC. It uses pattern matching to extract the coordinate field from the AtomC constructor. If the atom is an AtomC, it uses record syntax to access the x field of the coordinate and returns it wrapped in Just. If the atom is an AtomS, it returns Nothing since AtomS does not have a coordinate field.
@@
Listing 6: Retrieving the x-coordinate of an atom
@@ 3.1 The Molecule ADT
The Molecule data type is shown in Listing 7:
1data Molecule = Molecule {
@@
5
6data Atom = Atom {
@@
12
13data BondType = Bond {delocNum :: Integer, atomIDs :: Maybe [Integer]}
@@
18
19data ElementAttributes = ElementAttributes
@@
24
25type EquilibriumBondLength = Angstrom
26newtype Angstrom = Angstrom Double deriving (Read, Show, Eq)
Listing 7: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent Bonds (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over).
@@ 3.1.1 Example of a Molecule : Benzene
1benzene :: Molecule
2benzene = Molecule { atoms = [Atom i (elementAttributes e) (Coordinate x y z)
@@
5
6atomsData :: [(Int, Element, Double, Double, Double)]
7atomsData = [ (1, C, 0.0, 1.3948, 0.0), (2, C, 1.2079, 0.6974, 0.0), (3, C, 1.2079, -0.6974, 0.0),
8 (4, C, 0.0, -1.3948, 0.0), (5, C, -1.2079, -0.6974, 0.0), (6, C, -1.2079, 0.6974, 0.0),
9 (7, H, 0.0, 2.4732, 0.0), (8, H, 2.1431, 1.2366, 0.0), (9, H, 2.1431, -1.2366, 0.0),
10 (10, H, 0.0, -2.4732, 0.0), (11, H, -2.1431, -1.2366, 0.0), (12, H, -2.1431, 1.2366, 0.0) ]
11
12bondsData :: [((Int, Int), Bond)]
@@
14bondsData = [ ((1,2), single), ((1,6), single), ((2,3), single),
@@
Listing 8: A representation of the molecule benzene expressed in the Molecule data type, written as a .hs file, and using list comprehension to ensure readability, excluding the Shells currently.
@@
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as Bond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6].
@@ 3.2 The Orbital ADT
The representation captures quantum mechanical properties of atoms and can be instantiated such that it adheres to the principles governing the arrangement of electrons in an atom. Each shell is represented as a distinct entity, characterized by its principal quantum number. This aligns with the quantum mechanical description of atomic structure, where electrons occupy discrete energy levels or shells. By organizing the electronic configuration into separate shells, the representation accurately reflects the hierarchical nature of atomic orbitals and the increasing energy and complexity of the electronic structure as the principal quantum number increases.
@@
The use of a list to represent the shells implies that there could be an arbitrarily large finite number of shells (or even infinitely many), which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
@@
Within each shell, there are subshells ($s$, $p$, $d$, and $f$), which correspond to the different types of atomic orbitals. We do not include $g$ orbitals due to their rarity. Any interface to this data structure must respect the maximum number of electrons that can occupy each subshell, with $s$ subshells containing up to two electrons, $p$ subshells up to six electrons, $d$ subshells up to ten electrons, and $f$ subshells up to fourteen electrons. However, since the current implementation does not utilize dependent types, the correctness of the number of electrons in each subshell cannot be guaranteed at compile time and is left as future work. The responsibility of ensuring the proper electronic configuration lies with the user of the ADT.
Real atoms fill orbitals according to the interplay of $n$ and $l$ (the Aufbau principle), but well-known exceptions exist (e.g., chromium and copper) that reflect subtler electron–electron interactions. Spin is also relevant: each orbital can hold two electrons due to spin ±$\pm\frac{1}{2}$, yet here the ADT simply stores a count instead of tracking each electron’s spin. Finally, the $f$-orbital enumerations (e.g., Fxxx) are reasonable labels for real combinations of eigenstates, though in practice chemists often use different conventions or simply label them by magnetic quantum numbers.
Orbital hybridisation is also explicitly supported in this representation through the inclusion of the hybridComponents field in the Orbital data type. This feature not only allows for the accurate modeling of hybrid orbitals (such as sp^3, sp^2, etc.) but also enables predictions of chemical properties like acidity, as hybridisation can influence the electron density distribution and thus the reactivity of atoms in molecules.
Each atomic orbital contains the subshell type, electron count, and spatial orientation of the orbital. These properties can be used to construct the actual basis functions, but they are not the basis functions themselves. Furthermore, the representation allows for the specification of the spatial orientation of each orbital within a subshell through the use of the Coordinate type. This information is crucial for describing the shape and symmetry of atomic orbitals, particularly for $p$, $d$, and $f$ orbitals that have distinct orientations. However, it is important to note that the actual spatial representation of the orbitals would require the use of appropriate mathematical functions, such as Slater-type orbitals (STOs), spherical harmonics, or Gaussian functions. Slater-type orbitals are commonly used in quantum chemistry as they provide a more accurate representation of the electron density near the nucleus compared to Gaussian functions. However, these mathematical functions are not explicitly included in this ADT system, which focuses on the symbolic representation of the electronic configuration.
By designing the appropriate interface functions that operate on the Shells type, one can ensure that rules such as the Aufbau principle, Hund’s rule, and the Pauli exclusion principle are properly enforced when manipulating the electronic configuration. These interface functions would act as a layer of abstraction, encapsulating the logic and rules governing the distribution of electrons within the atom. By providing a well-defined set of operations that adhere to these principles, the interface can maintain the integrity and consistency of the electronic configuration, thereby accurately representing the charge of the molecule. The interface functions would handle the creation, modification, and querying of the Shells data structure, ensuring that any changes made to the electronic configuration are valid and compliant with quantum mechanical principles. This approach allows for a clear separation of concerns, where the Shells type focuses on data representation while the interface functions handle the logic and enforcement of the principles governing the electronic configuration.
The Aufbau principle can be enforced by ensuring that electrons are added to orbitals in the correct order of increasing energy. Hund’s rule can be implemented by maximizing the number of unpaired electrons in degenerate orbitals before pairing them. The Pauli exclusion principle can be enforced by limiting the maximum number of electrons in each orbital to two, with opposite spins (which could be implemented given more time). The use of type information can greatly assist in enforcing these principles and ensuring the validity of the electronic configuration. For example, the SubShell data type could be parameterized by the maximum number of orbitals allowed for each subshell type. This type-level information would prevent the creation of invalid subshells with more orbitals than permitted by quantum mechanical rules. Similarly, the electronCount field of the Orbital data type can be constrained to a range of values (e.g., 0 to 2) using type-level naturals or refined types. This ensures that the Pauli exclusion principle is upheld and no orbital contains more than two electrons.
1-- Basic orbital types
@@
3
4data P = Px | Py | Pz deriving (Show, Eq, Read)
5
6data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read)
7
8data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
9
10-- A type to wrap one of the pure orbital types,
11-- useful when describing the components of a hybrid orbital.
12data PureOrbital = PureSo So
13 | PureP P
14 | PureD D
15 | PureF F
16 deriving (Show, Eq, Read)
17
18-- The Orbital type includes an extra field for hybrid components.
19-- For a pure orbital, hybridComponents is Nothing.
20data Orbital subshellType = Orbital
@@
26
27-- A SubShell is a list of Orbitals all having the same subshell type.
28newtype SubShell subshellType = SubShell
@@
31
32-- A Shell has a principal quantum number and may contain s, p, d, and f subshells.
33data Shell = Shell
@@
40
41-- A Shells type is just a list of Shell.
42type Shells = [Shell]
@@
Listing 9: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells. Note some of the types here are not in blue, for simplicity.
@@
However, the current design does not explicitly account for the spin of electrons or the coupling between orbital and spin angular momenta, which play a significant role in determining the detailed electronic structure of atoms, particularly for elements with partially filled subshells.
@@
Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbital with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
@@
As mentioned, in Haskell, algebraic data types (ADTs) are constructed using sum and product type constructors, analogous to addition and multiplication in algebra (see Type Arithmetic). A sum type is a choice between different types (denoted by the | symbol), akin to an or; for example, a Maybe type represents a value that may be Nothing (like zero in addition) or Just a where a is any type. This is reflective of a disjoint sum in algebra, where a value can be one thing or another, but not both simultaneously. This allows the grammar to distinguish between empty values and zero values. As will be discussed, the orientation field of the Orbital data type is wrapped in a Maybe type, which naturally expresses whether an orbital has a specific orientation or not. For $s$ orbitals, the orientation is Nothing because $s$ orbitals are spherically symmetric and have no spatial orientation. In contrast, $p$ orbitals have specific orientations, and their orientation can be represented as Just Coordinate, where Coordinate is a type that can represent orientation in the $x$, $y$, and $z$ axes. On the other hand, a product type combines several values into one composite type, similar to how factors multiply to a product; a tuple (a, b) is an instance of a product type containing both a and b.
Furthermore, if the orbitals are not known, one can use the keyword undefined if they are unknown. Whilst risky, one could also implement an Unknown data type, which would be future work.